# S4ECG: Exploring the impact of long-range interactions for arrhythmia prediction

### v1

@@ Abstract
The electrocardiogram (ECG) exemplifies biosignal-based time series with continuous, temporally ordered structure reflecting cardiac physiological and pathophysiological dynamics. Detailed analysis of these dynamics has proven challenging, as conventional methods capture either global trends or local waveform features but rarely their simultaneous interplay at high temporal resolution. To bridge global and local signal analysis, we introduce S4ECG, a novel deep learning architecture leveraging structured state space models for multi-epoch arrhythmia classification. Our joint multi-epoch predictions significantly outperform single-epoch approaches by 1.0-11.6% in macro-AUROC, with atrial fibrillation specificity improving from 0.718-0.979 to 0.967-0.998, demonstrating superior performance in-distribution and enhanced out-of-distribution robustness. Systematic investigation reveals optimal temporal dependency windows spanning 10-20 minutes for peak performance. This work contributes to a paradigm shift toward temporally-aware arrhythmia detection algorithms, opening new possibilities for ECG interpretation, in particular for complex arrhythmias like atrial fibrillation and atrial flutter.
@@
Decision support systems , Electrocardiography , Time series analysis , Structured state space models , Long-range dependencies , Deep learning
journal: Computers in Biology and Medicine
[…]
organization=Carl von Ossietzky Universität Oldenburg,addressline=Ammerlaender Heerstr. 114-118, city=Oldenburg, postcode=26129, state=Lower Saxony, country=Germany
[…]
organization=German Heart Center of the Charité-University Medicine,addressline=Augustenburger Platz 1, city=Berlin, postcode=13353, state=Berlin, country=Germany
[…]
organization=Carl von Ossietzky Universität Oldenburg,addressline=Ammerlaender Heerstr. 114-118, city=Oldenburg, postcode=26129, state=Lower Saxony, country=Germany
@@ 1 Introduction
Clinical burden of cardiac arrhythmias Cardiovascular diseases remain the leading cause of mortality worldwide, with arrhythmias representing a significant subset of these conditions that can lead to sudden cardiac death, stroke, and heart failure if left undetected and untreated [1, 2, 3]. The clinical landscape is experiencing a notable shift toward atrial fibrillation (AF) as the most prevalent sustained arrhythmia, affecting millions of patients globally and imposing substantial healthcare burdens [4]. Early and accurate detection of arrhythmias is crucial for timely intervention [1, 3], with continuous electrocardiographic monitoring playing an increasingly vital role in modern cardiology practice [2].
@@
Challenges in AF detection The advent of portable devices and remote monitoring technologies has revolutionized arrhythmia detection [3, 5], enabling long-term continuous monitoring outside traditional clinical settings. However, the vast amounts of data generated by these devices present significant challenges for manual interpretation, creating an urgent need for automated algorithms that can reliably and accurately identify arrhythmic episodes. A particular challenge for arrhythmia detection algorithms remains the high rate of false positive alarms, accounting, for example, for almost 60% of the overall remote transmissions from implantable loop recorders [6]. While substantial progress has been made in automated ECG analysis research, enhancing model performance remains a pressing issue, particularly given the temporal complexity and variability inherent in cardiac rhythm disturbances that unfold over extended time periods.
@@
Long-range correlations The cardiovascular system exhibits well-documented long-range temporal correlations, particularly evident in heart rate variability patterns during different physiological states [7]. These long-range interactions manifest across multiple timescales, from beat-to-beat variations to circadian rhythms, and have been shown to carry diagnostic information for cardiac pathology detection [8]. For instance, healthy heart dynamics exhibit multifractal complexity persisting for at least 700 beats (approximately 10 minutes) - a hallmark of physiological control markedly reduced in cardiac pathology [9]. Similarly, detrended fluctuation analysis of 24-hour heartbeat recordings reveals that the characteristic scale-invariant long-range correlations in healthy subjects persist across scales from $10^{2}$ to $10^{3}$ beats (approximately 1 to 20 minutes), whereas pathologic dynamics deviate from this behavior at these scales [10]. Additionally, cardiac electrophysiology and autonomic tone exhibit pronounced circadian rhythms, and arrhythmic events cluster by time of day [11]. The presence of such temporal dependencies suggests that arrhythmia detection algorithms could benefit substantially from incorporating extended temporal context.
@@
Algorithmic approaches Automated ECG analysis has been an active area of research for several decades, with traditional approaches focusing primarily on handcrafted feature extraction and classical machine learning algorithms. Early methods relied on morphological features, frequency domain characteristics, and statistical measures derived from beat-to-beat intervals [12]. These approaches, while providing interpretable results, often struggled with the variability inherent in real-world ECG recordings and required extensive domain expertise for feature engineering. The development of deep learning has transformed ECG analysis, with convolutional neural networks (CNNs) and (to a lesser extent) long short-term memory (LSTM) networks, emerging as the dominant paradigm for automated arrhythmia detection [12].
@@
Shortcomings of existing approaches However, most existing approaches have primarily focused on single-epoch classification, where individual 5-30 second segments are analyzed independently to identify the underlying cardiac rhythm [12]. In line with conventions in the sleep staging literature, we refer to these segments as epochs, while recognizing the potential for confusion with training epochs commonly used in the machine learning domain. This paradigm, while computationally efficient and conceptually straightforward, inherently limits the temporal context available for decision-making. Single-epoch models cannot capture rhythm transitions, paroxysmal episodes, or gradual changes in cardiac rhythm that may unfold over minutes to hours [3]—temporal patterns that are clinically significant for accurate arrhythmia characterization. Single-epoch models tend to misjudge contiguous arrhythmic events by making inconsistent predictions across adjacent segments, leading to inappropriate rhythm segmentation boundaries and fragmented episode detection. This segmentation error stems from the lack of temporal overview that would enable recognition of sustained arrhythmic patterns extending beyond individual windows. Furthermore, most existing ECG analysis architectures are fundamentally constrained by their reliance on traditional deep learning components, primarily CNNs and LSTM [12]. While CNNs excel at capturing local morphological features within individual heartbeats, they are inherently limited in modeling long-range temporal dependencies that extend beyond their receptive field. LSTMs, despite their theoretical ability to capture sequential dependencies, suffer from vanishing gradient problems and computational inefficiency when processing extended sequences, making them impractical for multi-epoch analysis spanning tens of minutes. These architectural limitations have prevented the field from leveraging recent advances in sequence modeling, particularly structured state space models, which offer superior capabilities for efficient long-range dependency capture.
@@
Our primary contribution lies in adapting and extending the encoder-predictor paradigm from sleep staging [13] to ECG analysis, employing structured state space models (S4) [14] at both epoch-level and sequence-level processing stages to enable efficient capture of long-range dependencies. Firstly, we confirm the superiority of S4-based encoders over widely used CNN-based encoders in line with prior work [15]. Secondly and most importantly, we demonstrate that multi-epoch models consistently outperform single-epoch approaches across diverse datasets and evaluation scenarios.
We present the first comprehensive study to systematically investigate the optimal temporal window for ECG arrhythmia detection, evaluating model performance across temporal contexts ranging from 2 to 60 epochs, i.e., 1 to 30 minutes of continuous ECG data at an epoch length of 30 seconds. Our findings reveal consistent optimal performance in the 20-40 epoch range, i.e., 10 to 20 minutes, suggesting fundamental characteristics of cardiac rhythm analysis that extend beyond dataset-specific artifacts. Our evaluation encompasses training on large-scale and medium-sized datasets, followed by rigorous out-of-distribution testing on medium-sized and smaller benchmark databases from PhysioNet [16]. This comprehensive evaluation framework spans diverse acquisition protocols, patient populations, and clinical contexts, enabling thorough assessment of model robustness and generalizability across real-world deployment scenarios. The consistency of our findings across these diverse datasets provides strong evidence for the generalizability of multi-epoch approaches in clinical ECG analysis. To summarize, we put forward the following technical contributions:
@@
    We assess hierarchical prediction models leveraging structured state space models that showed outstanding performance in sleep stage prediction [13] for the purpose of arrhythmia prediction. The proposed models, in particular when trained on large-scale datasets such as Icentia11k [17], show strong predictive performance, also when evaluated on out-of-distribution data.
@@
    We provide robust evidence for the advantages of jointly predicting multiple prediction epochs at once as opposed to a single epoch at a time, as predominantly considered in the literature, both in terms of in-distribution and out-of-distribution performance. We present qualitative evidence for the advantages of such multi-epoch prediction models.
@@ 2.1 Datasets and experimental setup
Datasets This study investigates the effectiveness of multi-epoch deep learning models for arrhythmia detection using a comprehensive evaluation framework encompassing training, in-distribution (ID), and out-of-distribution (OOD) assessments. We use two datasets for model training and ID evaluation, and—for OOD evaluation—combine one of these training datasets with two additional external datasets to assess robustness and generalizability. All datasets are publicly accessible through PhysioNet [16], ensuring reproducibility. Table 1 and 2 provide an overview of the four datasets, including patient counts, recording durations, sampling frequencies, and rhythm-type distributions. The substantial differences in dataset scale, temporal resolution, and class prevalence enable a thorough evaluation of multi-epoch performance under varying conditions, from ID settings that mirror training characteristics to challenging OOD scenarios. Detailed dataset and preprocessing descriptions are provided in Section A.
@@
Icentia11k[17] Training, ID eval 11,000 250 Variable (hours-weeks) $\sim$110,000
LTAFDB[18] Training, ID/OOD eval 84 128 24-25 hours $\sim$2,000
AFDB[19] OOD eval 25 250 $\sim$10 hours $\sim$250
MITDB[20] OOD eval 47 360 30 minutes $\sim$24
Experimental setup We convert rhythm annotation timestamps into relative fractions of the considered rhythm types per prediction epoch, see A. We use these fractional labels as prediction targets using a binary crossentropy loss function optimized through an AdamW optimizer. The macro-averaged area under the receiver operating characteristic curve (AUROC) serves as primary metric. For clinical interpretability, we also report AF detection specificity at a fixed sensitivity of 0.9, consistent with the performance levels achieved by FDA-cleared wearable AF detection devices [21, 22]. We refer the reader to B for extensive details on the experimental setup.
Table 2: Rhythm-type distribution across datasets (% of epochs).
@@ 2.2 S4ECG architecture
This work builds on prior advances in sleep staging [13], which conducted extensive architecture searches for long time-series classification. It provided evidence for the superiority of S4 layers over LSTMs or transformer architectures in a closely related setting. We therefore refrain from excessive ablation studies and focus exclusively on the best-performing model identified in prior work. The S4ECG model implements an encoder-predictor paradigm that leverages structured state space (S4) layers [14] as core components across two processing stages.
@@
Architecture overview The model processes multi-epoch ECG sequences through a hierarchical encoder-predictor design. An input sequence of length $L$ samples is first segmented into $N$ non-overlapping epochs of fixed length $L_{\text{epoch}}=3840$ samples (30 seconds at 128 Hz), where $N=L/L_{\text{epoch}}$. Each epoch is encoded independently; the resulting token sequence is then modeled by a predictor to capture inter-epoch dependencies. A classification head produces per-epoch rhythm predictions.
@@
Epoch-level encoder Each 30-second epoch is passed through a convolutional front-end (two 1D convolutional layers with 128 channels, kernel size 3, and stride 2), which reduces the temporal dimension from 3840 to 960 samples, followed by an S4 stack (model dimension 512, state dimension 64, 4 layers, bidirectional). A pooling operation compresses each epoch to a single 512-dimensional token.
@@
Multi-epoch predictor The sequence of epoch tokens is processed by a second, four-layer S4 module (model dimension 512, bidirectional) that captures long-range temporal dependencies across epochs. The output is fed to a linear classification head to produce rhythm predictions.
@@
Multi-epoch vs. single-epoch comparison Unlike conventional single-epoch models that process fixed 30-second segments (input size = 3,840 samples), as shown in Figure 1(a), our S4ECG model shown in Figure 1(b) processes variable-length sequences containing multiple epochs. For example, with input size 38,400, the model processes $N=10$ epochs spanning 5 minutes of continuous ECG. This enables modeling of patterns extending beyond individual segments, such as paroxysmal episodes and rhythm transitions.
@@
The inclusion of the multi-epoch predictor and the formulation of the task as a joint prediction over several epochs constitute the central novelty of our approach. We systematically compare this against conventional single-epoch models across temporal contexts. For LTAFDB, we evaluate 2 to 60 epochs (1 to 30 minutes of ECG), whereas for the large-scale Icentia11k dataset, computational constraints limit evaluation to 10 to 60 epochs (5 to 30 minutes).
(a) Single-epoch architecture for baseline comparison.
@@
(b) Multi-epoch S4ECG architecture with hierarchical encoder-predictor design
@@
Figure 1: Architecture comparison: (a) single-epoch baseline model and (b) multi-epoch S4ECG model with temporal dependency modeling.
@@
Single-epoch baselines We consider two baseline models that operate on a single epoch as input. On the one hand, we consider a xResNet1d50 model [23] as representative for the predominantly used CNNs for this task. On the other hand, we consider a S4-based single-epoch baseline model, which emerged as strongest single-epoch backbone in [13].

### v2

# Multi-Window Temporal Analysis for Enhanced Arrhythmia Classification: Leveraging Long-Range Dependencies in Electrocardiogram Signals

@@ Abstract
Objective. Arrhythmia classification from electrocardiograms (ECGs) suffers from high false positive rates and limited cross-dataset generalization, particularly for atrial fibrillation (AF) detection where specificity ranges from 0.72 to 0.98 using conventional 30-second analysis windows. While conventional deep learning approaches analyze isolated 30-second ECG windows, many arrhythmias, particularly atrial fibrillation and atrial flutter, exhibit diagnostic features that emerge over extended time scales. Approach. We introduce S4ECG, a deep learning architecture based on structured state-space models (S4), designed to capture long-range temporal dependencies by jointly analyzing multiple consecutive ECG windows spanning up to 20 minutes. We evaluated S4ECG on four publicly available databases for multi-class arrhythmia classification, including systematic cross-dataset evaluations to assess out-of-distribution robustness. Main results. Multi-window analysis consistently outperformed single-window approaches across all datasets, improving the macro-averaged area under the receiver operating characteristic curve (AUROC) by 1.0-11.6 percentage points. For AF detection specifically, specificity increased from 0.718-0.979 (single-window) to 0.967-0.998 (multi-window) at a fixed sensitivity threshold, representing a 3-10 fold reduction in false positive rates. Significance. Comparative analysis against convolutional neural network baselines demonstrated superior performance of the S4 architecture. Cross-dataset evaluation revealed that multi-window approaches substantially improved generalization performance, with smaller performance degradation when models were tested on held-out datasets from different institutions and acquisition protocols. A systematic investigation revealed optimal diagnostic windows of 10-20 minutes, beyond which performance plateaus or degrades. These findings demonstrate that structured incorporation of extended temporal context enhances both arrhythmia classification accuracy and cross-dataset robustness. The identified optimal temporal windows provide practical guidance for ECG monitoring system design and may reflect underlying physiological timescales of arrhythmogenic dynamics.
@@
Decision support systems, Electrocardiography, Arrhythmia, Atrial Fibrillation, Time series analysis, Deep learning
articletype: Paper
@@ 1 Introduction
Clinical burden of cardiac arrhythmias Cardiovascular diseases remain the leading cause of mortality worldwide, with arrhythmias representing a significant subset of these conditions that can lead to sudden cardiac death, stroke, and heart failure if left undetected and untreated (Ansari et al., 2023; Svennberg et al., 2025; Ko et al., 2025). The clinical landscape is experiencing a notable shift toward atrial fibrillation (AF) as the most prevalent sustained arrhythmia, affecting millions of patients globally and imposing substantial healthcare burdens (Conrad et al., 2025). Early and accurate detection of arrhythmias is crucial for timely intervention (Ansari et al., 2023; Ko et al., 2025), with continuous electrocardiographic monitoring playing an increasingly vital role in modern cardiology practice (Svennberg et al., 2025).
@@
Challenges in AF detection The advent of portable devices and remote monitoring technologies has revolutionized arrhythmia detection (Ko et al., 2025; Abdelrazik et al., 2025), enabling long-term continuous monitoring outside traditional clinical settings. However, the vast amounts of data generated by these devices present significant challenges for manual interpretation, creating an urgent need for automated algorithms that can reliably and accurately identify arrhythmic episodes. A particular challenge for arrhythmia detection algorithms is the high rate of false positive alarms, accounting, for example, for almost 60% of the overall remote transmissions from implantable loop recorders (Covino and Russo, 2024). While substantial progress has been made in automated ECG analysis research, enhancing model performance remains a pressing issue, particularly given the temporal complexity and variability inherent in cardiac rhythm disturbances that unfold over extended time periods.
@@
Long-range correlations The cardiovascular system exhibits well-documented long-range temporal correlations, particularly evident in heart rate variability patterns during different physiological states (Bunde et al., 2000). These long-range interactions manifest across multiple timescales, from beat-to-beat variations to circadian rhythms, and have been shown to carry diagnostic information for cardiac pathology detection (Agliari et al., 2020). For instance, healthy heart dynamics exhibit multifractal complexity persisting for at least 700 beats (approximately 10 minutes) - a hallmark of physiological control markedly reduced in cardiac pathology (Ivanov et al., 1999). Similarly, detrended fluctuation analysis of 24-hour heartbeat recordings reveals that the characteristic scale-invariant long-range correlations in healthy subjects persist across scales from $10^{2}$ to $10^{3}$ beats (approximately 1 to 20 minutes), whereas pathologic dynamics deviate from this behavior at these scales (Peng et al., 1995). Additionally, cardiac electrophysiology and autonomic tone exhibit pronounced circadian rhythms, and arrhythmic events cluster by time of day (Kelters et al., 2025). The presence of such temporal dependencies suggests that arrhythmia detection algorithms could benefit substantially from incorporating extended temporal context.
@@
Algorithmic approaches Automated ECG analysis has been an active area of research for several decades, with traditional approaches focusing primarily on handcrafted feature extraction and classical machine learning algorithms. Early methods relied on morphological features, frequency domain characteristics, and statistical measures derived from beat-to-beat intervals (Murat et al., 2021). These approaches, while providing interpretable results, often struggled with the variability inherent in real-world ECG recordings and required extensive domain expertise for feature engineering. The development of deep learning has transformed ECG analysis, with convolutional neural networks (CNNs) and (to a lesser extent) long short-term memory (LSTM) networks, emerging as the dominant paradigm for automated arrhythmia detection (Murat et al., 2021).
@@
Shortcomings of existing approaches However, most existing approaches have primarily focused on single-window classification, where individual 5-30 second windows are analyzed independently to identify the underlying cardiac rhythm (Murat et al., 2021). This paradigm, while computationally efficient and conceptually straightforward, inherently limits the temporal context available for decision-making. Single-window models cannot capture rhythm transitions, paroxysmal episodes, or gradual changes in cardiac rhythm that may unfold over minutes to hours (Ko et al., 2025)—temporal patterns that are clinically significant for accurate arrhythmia characterization. Single-window models tend to misjudge contiguous arrhythmic events by making inconsistent predictions across adjacent segments, leading to inappropriate rhythm segmentation boundaries and fragmented episode detection. This segmentation error stems from the lack of temporal overview that would enable recognition of sustained arrhythmic patterns extending beyond individual windows. Furthermore, most existing ECG analysis architectures are fundamentally constrained by their reliance on traditional deep learning components, primarily CNNs and LSTM (Murat et al., 2021). While CNNs excel at capturing local morphological features within individual heartbeats, they are inherently limited in modeling long-range temporal dependencies that extend beyond their receptive field. LSTMs, despite their theoretical ability to capture sequential dependencies, suffer from vanishing gradient problems and computational inefficiency when processing extended sequences, making them impractical for multi-window analysis spanning tens of minutes. These architectural limitations have prevented the field from leveraging recent advances in sequence modeling, particularly structured state space models (S4) (Gu et al., 2022) , which offer superior capabilities for efficient long-range dependency capture.
@@
Our primary contribution lies in adapting and extending the encoder-predictor paradigm from sleep staging (Wang and Strodthoff, 2025a) to ECG analysis, employing S4 models at both window-level and sequence-level processing stages to enable efficient capture of long-range dependencies. Firstly, we confirm the superiority of S4-based encoders over widely used CNN-based encoders in line with prior work (Al-Masud et al., 2025). Secondly and most importantly, we demonstrate that multi-window models consistently outperform single-window approaches across diverse datasets and evaluation scenarios.
We present the first comprehensive study to systematically investigate the optimal temporal window for ECG arrhythmia detection, evaluating model performance across temporal contexts ranging from 2 to 60 windows, i.e., 1 to 30 minutes of continuous ECG data at a window length of 30 seconds. We emphasize that we work with a fixed window size of 30s throughout this work, longer temporal context is implemented by considering multiple consecutive windows at once. Our findings reveal consistent optimal performance in the 20-40 window range, i.e., 10 to 20 minutes, suggesting fundamental characteristics of cardiac rhythm analysis that extend beyond dataset-specific artifacts. Our evaluation encompasses training on large-scale and medium-sized datasets, followed by rigorous out-of-distribution testing on medium-sized and smaller benchmark databases from PhysioNet (Goldberger et al., 2000). This comprehensive evaluation framework spans diverse acquisition protocols, patient populations, and clinical contexts, enabling thorough assessment of model robustness and generalizability across real-world deployment scenarios. The consistency of our findings across these diverse datasets provides strong evidence for the generalizability of multi-window approaches in clinical ECG analysis. To summarize, we put forward the following technical contributions:
@@
    We assess hierarchical prediction models leveraging structured state space models that showed outstanding performance in sleep stage prediction (Wang and Strodthoff, 2025a) for the purpose of arrhythmia prediction. The proposed models, in particular when trained on large-scale datasets such as Icentia11k (Tan et al., 2022), show strong predictive performance, also when evaluated on out-of-distribution data.
@@
    We provide robust evidence for the advantages of jointly predicting multiple prediction windows at once as opposed to a single window at a time, as predominantly considered in the literature, both in terms of in-distribution and out-of-distribution performance. We present qualitative evidence for the advantages of such multi-window prediction models.
@@ 2.1 Datasets and experimental setup
Datasets This study investigates the effectiveness of multi-window deep learning models for arrhythmia detection using a comprehensive evaluation framework encompassing training, in-distribution (ID), and out-of-distribution (OOD) assessments. We use two datasets for model training and ID evaluation, and—for OOD evaluation—combine one of these training datasets with two additional external datasets to assess robustness and generalizability. All datasets are publicly accessible through PhysioNet (Goldberger et al., 2000), ensuring reproducibility. Table 1 and 2 provide an overview of the four datasets, including patient counts, recording durations, sampling frequencies, and rhythm-type distributions. The substantial differences in dataset scale, temporal resolution, and class prevalence enable a thorough evaluation of multi-window performance under varying conditions, from ID settings that mirror training characteristics to challenging OOD scenarios. We emphasize that all deep learning models were trained directly on the raw ECG time-series signals, without applying additional preprocessing steps such as baseline-wander correction or signal filtering. This design choice was intentional: it enables the model to learn directly from raw signals, thereby improving generalizability across different acquisition systems and avoiding potential information loss from aggressive filtering or hand-crafted preprocessing pipelines. Detailed dataset and preprocessing descriptions are provided in Section LABEL:app:datasets.
@@
Icentia11k(Tan et al., 2022) Training, ID eval 11,000 250 Variable (hours-weeks) $\sim$110,000
LTAFDB(Petrutiu et al., 2006) Training, ID/OOD eval 84 128 24-25 hours $\sim$2,000
AFDB(Moody and Mark, 2006) OOD eval 25 250 $\sim$10 hours $\sim$250
MITDB(Moody and Mark, 2001) OOD eval 47 360 30 minutes $\sim$24
Experimental setup We convert rhythm annotation timestamps into relative fractions of the considered rhythm types per prediction window, see Section LABEL:app:datasets. We use these fractional labels as prediction targets using a binary crossentropy loss function optimized through an AdamW optimizer. The macro-averaged area under the receiver operating characteristic curve (AUROC) serves as primary metric. For clinical interpretability, we also report AF detection specificity at a fixed sensitivity of 0.9, consistent with the performance levels achieved by FDA-cleared wearable AF detection devices (Perez et al., 2019; Bumgarner et al., 2018). We refer the reader to Section LABEL:app:setup for extensive details on the experimental setup.
Table 2: Rhythm-type distribution across datasets (% of windows).
@@ 2.2 S4ECG architecture
This work builds on prior advances in sleep staging (Wang and Strodthoff, 2025a), which conducted extensive architecture searches for long time-series classification. It provided evidence for the superiority of S4 layers over LSTMs or transformer architectures in a closely related setting. We therefore refrain from excessive ablation studies and focus exclusively on the best-performing model identified in prior work. The S4ECG model implements an encoder-predictor paradigm that leverages S4 layers as core components across two processing stages.
@@
Architecture overview The model processes multi-window ECG sequences through a hierarchical encoder-predictor design. An input sequence of length $L$ samples is first segmented into $N$ non-overlapping windows of fixed length $L_{\text{window}}=3840$ samples (30 seconds at 128 Hz), where $N=L/L_{\text{window}}$. Each window is encoded independently; the resulting token sequence is then modeled by a predictor to capture inter-window dependencies. A classification head produces per-window rhythm predictions. Detailed architectural and training hyperparameters are provided in Section LABEL:app:repro.
@@
Window-level encoder Each 30-second window is passed through a convolutional front-end (two 1D convolutional layers with 128 channels, kernel size 3, and stride 2), which reduces the temporal dimension from 3840 to 960 samples, followed by an S4 stack (model dimension 512, state dimension 64, 4 layers, bidirectional). A pooling operation compresses each window to a single 512-dimensional token.
@@
Multi-window predictor The sequence of window tokens is processed by a second, four-layer S4 module (model dimension 512, bidirectional) that captures long-range temporal dependencies across windows. The output is fed to a linear classification head to produce rhythm predictions.
@@
Multi-window vs. single-window comparison Unlike conventional single-window models that process fixed 30-second segments (input size = 3,840 samples), as shown in Figure 1(a), our S4ECG model shown in Figure 1(b) processes variable-length sequences containing multiple windows. For example, with input size 38,400, the model processes $N=10$ windows spanning 5 minutes of continuous ECG. This enables modeling of patterns extending beyond individual segments, such as paroxysmal episodes and rhythm transitions.
@@
The inclusion of the multi-window predictor and the formulation of the task as a joint prediction over several windows constitute the central novelty of our approach. We systematically compare this against conventional single-window models across temporal contexts. For LTAFDB, we evaluate 2 to 60 windows (1 to 30 minutes of ECG); for the large-scale Icentia11k dataset, we narrow this to 10 to 60 windows (5 to 30 minutes) for computational efficiency.
(a) Single-window architecture for baseline comparison.
@@
(b) Multi-window S4ECG architecture with hierarchical encoder-predictor design
@@
Figure 1: Architecture comparison: (a) single-window baseline model and (b) multi-window S4ECG model with temporal dependency modeling.
@@
Single-window baselines We consider two baseline models that operate on a single window as input. On the one hand, we consider a xResNet1d50 model (Strodthoff et al., 2020) as representative for the predominantly used CNNs for this task. On the other hand, we consider a S4-based single-window baseline model, which emerged as strongest single-window backbone in (Wang and Strodthoff, 2025a).